OpenStudy (anonymous):

If x varies inversely as y and directly as t, and x = 12 when t = 10 and y = 25, find y when x is 6 and t = 3.

OpenStudy (mokeira):

\[x=k\frac{ t }{ y }\] \[12=k \frac{ 10 }{ 25 }\] k=30 \[6=30(\frac{ 3 }{y })\] \[6=\frac{ 90 }{y }\] \[y=\frac{ 90 }{6 }\] simplify to get y hope you understand my steps